OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -2) and a directrix at y = 2.

OpenStudy (anonymous):

@Hero

hero (hero):

@Nairz, what are your thoughts regarding this problem? Do you have an approach for solving this?

OpenStudy (anonymous):

Im assuming the vertex is the origin since the directrix and the focus are both 2 units away from the origin along the same line.

OpenStudy (anonymous):

@Hero

hero (hero):

Consider this: If you have two points, the focus \((x_1, y_1)\) and the directrix \((x_2, y_2)\) then you can insert those points in to the following formula to find the standard form of the equation of a parabola: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\)

OpenStudy (anonymous):

And that will give me the equation to the parabola?

hero (hero):

Yes. Notice that in this case: Focus: \((x_1, y_1) = (0,-2)\) Directrix: \((x_2, y_2) = (x,2)\)

hero (hero):

You'll have to do a bit of simplification after inserting the points in order to express the equation in standard form.

OpenStudy (anonymous):

yes. Thanks!

hero (hero):

@Nairz, mind showing your work for this? Let's see what equation you come up with.

OpenStudy (anonymous):

ummm I got \[x=2\sqrt{2}\]

OpenStudy (anonymous):

ill show my work in a second

OpenStudy (anonymous):

\[\large (x-0)^2 + (y+2)^2 = (x-x)^2 + (y-2)^2\]

OpenStudy (anonymous):

\[\large x^2 + (y+2)^2 = (y+2)^2\]

OpenStudy (anonymous):

\[\large x^2 = (y-2)^2 - (y+2)^2\]

OpenStudy (anonymous):

\[\large x^2 = -8\]

OpenStudy (anonymous):

\[\large x=2\sqrt{2}\]

hero (hero):

Okay, hang on a second.

hero (hero):

\((y - 2)^2 - (y + 2)^2\) is actually "difference of squares". You should use the difference of squares formula to simplify that properly.

OpenStudy (anonymous):

i dont have a formula sheet in front of me

hero (hero):

Difference of squares formula: \(a^2 - b^2 = (a + b)(a - b)\)

hero (hero):

In this case a = y - 2 b = y + 2

OpenStudy (anonymous):

so its \[\large ((y-2)+ (y+2))((y-2)-(y+2))\]

hero (hero):

Don't forget the \(x^2\) equals part

OpenStudy (anonymous):

simplified: \[\large x^2 = 2y(-4)\]

OpenStudy (anonymous):

so x^2 = -8y

OpenStudy (anonymous):

which isnt an option...

hero (hero):

Finish isolating y

OpenStudy (anonymous):

y=x^2/-8

hero (hero):

Note that the equation can also be written in the form \(y = -\dfrac{1}{8}x^2\)

OpenStudy (anonymous):

Ok... I just realized I was looking at the wrong problem's answers... That is there. Thanks

OpenStudy (anonymous):

Can you work through another one of these with me? Just to make sure I do it correctly?